A hotel in Hong Kong has 4 room service maids to clean daily 31 standard rooms, 20 deluxe rooms and 9 suites. The standard times for servicing rooms are as follows: standard rooms 20 minutes, deluxe rooms 25 minutes and suites 40 minutes. All room service maids work an 7-hour day. Calculate the efficiency of the maids. (provide percentage value round to two decimal places).
Answer #1

Time taken to service 31 standard rooms = 20/60 = 0.33 = 31*0.33 = 10.23 hrs

Time taken to service 20 deluxe rooms = 25/60 = 0.41 = 20*0.41 = 8.20 hrs

Time taken to service 9 suites = 40/60 = 0.66 = 9*0.66 = 5.94 hrs

4 room service maids spending 7 Hours = 7*4 = 28 hours

Total time = 10.23 + 8.20 + 5.94 = 24.37

Efficiency = Standard Labor Hours / Amount of time worked *100 = 28/24.37 *100 = 114.89%